A car caught in a heavy traffic situation along rizal avenue moves at constant speed covering of 3.0 m in 5.0 s. (A) what is yhe speed of the car? (B) how far will the object move in 10.0s?
1
Expert's answer
2020-10-15T10:50:31-0400

a)


"v=\frac{s}{t}=\frac{3}{5}=0.6\frac{m}{s}"

b)


"d=vT=(0.6)(10)=6.0\ m"